nene23:

Sharon’s turtle escaped from her backyard sometime in the last few hours. According to her calculations, the farthest the turtle could have gone is 4 blocks down the road in either direction. If Sharon lives on the 112th block of town, which equation can be used to find the block numbers that represent the farthest distance that the turtle may be?

Mercury:

choices: |x – 112| = 0 |x – 112| = 4 |4 – 112| = x |112 – 4| = x the turtle is at most 4 blocks up or 4 blocks down in other words, the difference between the turtle and 112 must either be +4 or -4 depending on direction translate this idea into the appropriate absolute value equation